(13%) Problem 4: A charged particle (q-7.5 10-10 C) experiences a force of F-2.25i-6.2j N in an electric field 50% Part (a) Write an expression for the electric field vector E to which the charge is subject, in terms of the force F 50% Part b) Assume this field is generated by a point charge of Q-5x 109 C. How far away is this charge located? Give your answer in meters Grade Summa
Answer #1

Given F (2.252-6.2j )N

Use Formula F=Eq

F/q=E

10 Ε = (2.25i-62j)/(7.5 * 10- )

E (0.3i-0.8267 1010

Part A) ANSWER: E (Зі _ 8.267) * 109 Ус

==============================================

E=kq/r^{2}

E (Зі _ 8.267) * 109 Ус

E-8.7945 * 10%V/C

8.7945 10 NIC k5 10Cr

8.7945*10^{9}=9*10^{9}*5*10^{-9}/r^{2}

8.7945 10-95/r2

r^{2}=9*5/8.7945*10^{9}

7.2 45 * 10-9/8.7945

ANSWER: r=7.15*10^{-5}m

============================